OpenStudy (anonymous):

what is the slope intercept form of this standard intercept form? the equation is 2x+10y=36.

OpenStudy (anonymous):

for x=0 , y = 3.6 for y = 0 , x = 18

OpenStudy (anonymous):

10y=-2x+36 5y=-1+18 comparing with, y=mx+c m(slope)=-1